OpenStudy (anonymous):

Solve: 2(x + 2) − (x + 1) = 8

OpenStudy (anonymous):

2(x+2)-(x+1)= 8 distribute the 2 and the negative: 2x+4-x-1=8 combine terms: x+3=8 subtract the 3: x=8-3 x=5

OpenStudy (anonymous):

@ronocthede thankss!!!